Let V and W be n−dimensional vector spaces, and let T : V → W be a linear transformation.

Suppose β is a basis for V . Prove that T is an isomorphism if and only if T(β) is a basis for W.



1
Expert's answer
2021-07-21T07:26:24-0400

1) "T" is an isomorphism.

We need to show that "T(\beta)" is a basis for "W" .

The set "T(\beta)=\{T(\beta_1), …, T(\beta_n)\}" consists of "n" vectors. It suffices to show that "T(\beta_1), …, T(\beta_n)" are linearly independent.

Let "\sum \limits_{i=1}^na_iT(\beta_i)=0" for some "a_i" .

Then "\sum \limits_{i=1}^na_iT(\beta_i)= \sum \limits_{i=1}^n T(a_i\beta_i)=T\left(\sum\limits_{i=1}^na_i\beta_i
\right)=0"

Since "T" is an isomorphism, it follows that "Tv=0" only if "v=0" .

Therefore, "\sum\limits_{i=1}^na_i\beta_i=0" . Since "\beta" is a basis for "V" , all "a_i=0" .

Hence, "T(\beta)" is linearly independent.


2) "T(\beta)" is a basis for "W" .

We need to show that "T" is an isomorphism.


"T" is injective if and only if "\text{Ker}\ T=0" .

Suppose that "v\in \text{Ker} \ T" and "v=\sum\limits_{i=1}^na_i\beta_i" for some "a_i" .

"Tv=T\left(\sum\limits_{i=1}^na_i\beta_i
\right)=\sum\limits_{i=1}^nT(a_i\beta_i)=\sum\limits_{i=1}^na_iT(\beta_i)=0."

Since "T(\beta)" is a basis for "W" , it follows that all "a_i=0" . Then we have that "v=0" and "\text{Ker}\ T=0" .


"T" is surjective if and only if for all "w\in W" there exists "v\in V" such that "Tv=w" .

Let "w=\sum\limits_{i=1}^na_iT(\beta_i)" for some "a_i" .

Then "w=\sum\limits_{i=1}^na_iT(\beta_i)=\sum\limits_{i=1}^nT(a_i\beta_i)=T\left(
\sum\limits_{i=1}^na_i\beta_i
\right)" , where "\sum\limits_{i=1}^na_i\beta_i\in V."


"T" is injective and surjective. So, "T" is an isomorphism.